# Towards Principled Assessment of Tabular Data Synthesis Algorithms

### v1

@@ Abstract
Data synthesis has been advocated as an important approach for utilizing data while protecting data privacy. A large number of tabular data synthesis algorithms (which we call synthesizers) have been proposed. Some synthesizers satisfy Differential Privacy, while others aim to provide privacy in a heuristic fashion. A comprehensive understanding of the strengths and weaknesses of these synthesizers remains elusive due to lacking principled evaluation metrics and missing head-to-head comparisons of newly developed synthesizers that take advantage of diffusion models and large language models with state-of-the-art marginal-based synthesizers.
@@
In this paper, we present a principled and systematic evaluation framework for assessing tabular data synthesis algorithms. Specifically, we examine and critique existing evaluation metrics, and introduce a set of new metrics in terms of fidelity, privacy, and utility to address their limitations. Based on the proposed metrics, we also devise a unified objective for tuning, which can consistently improve the quality of synthetic data for all methods. We conducted extensive evaluations of 8 different types of synthesizers on 12 datasets and identified some interesting findings, which offer new directions for privacy-preserving data synthesis.
1 Introduction
@@
Data-driven decision-making has emerged as the prevailing approach to advance science, industrial applications, and governance, creating the necessity to share and publish tabular data. At the same time, growing concerns about the privacy breaches caused by data disclosure call for data publishing approaches that preserve privacy. One increasingly advocated and adopted approach to reduce privacy risks while sharing data is to release synthetic data. Ideally, synthetic data can effectively fit any data processing workflow designed for the original data without privacy concerns. Data synthesis initiatives have been promoted not only by the research community [58, 29] but also among non-profit organizations [45, 61] and government agencies [5, 43, 44].
@@
In this paper, we study data synthesis algorithms for tabular data, which we call synthesizers. In recent years, a plethora of synthesizers have been proposed, which can be roughly categorized into two groups: statistical and deep generative. Statistical synthesizers [69, 41, 39, 70] use low-order marginals of the real data to create synthetic datasets that match real distributions. They were the best-performing algorithms in NIST competitions [43, 44]. Deep generative synthesizers [48, 30, 65, 32, 7], on the other hand, learn the data distribution from real data records and generate synthetic instances by sampling from the distribution. With the recent development in deep generative models (e.g., diffusion models [26] and large language models [64, 54]), new synthesizers were proposed [32, 7, 31, 68] to extend these successes to the realm of tabular data synthesis.
@@
While recent state-of-the-art approaches achieve compelling results in synthesizing authentic data, a comprehensive understanding of the strengths and weaknesses of different synthesizers remains elusive. We are unaware of any head-to-head comparison of the best statistical synthesizers [41, 39, 70] against the best deep generative synthesizers [32, 7, 31].
@@
In addition, there is a lack of principled and widely accepted evaluation metrics. It is known that evaluating synthesizers is inherently difficult [59, 12, 2, 29], and qualitative evaluation of tabular data through visual inspection is also infeasible. Table 1 shows that existing papers use different evaluation metrics, resulting in substantially disparate conclusions.
@@
Table 1: Main evaluation protocols for synthesizers in the literature. “-” means that the synthesizer does not consider this property. “ ✓ ” denotes that the model is viewed as the state-of-the-art method by many related studies [40, 20, 46, 58]. “ ✓   ” denotes that the model is frequently used in recent papers[32, 34, 31] for comparison as a state-of-the-art method.
@@
PGM [39, 41] Differentially Private Probabilistic Graphical Model Total Variation Distance Range Query
@@
PrivSyn [70] Differentially Private Non-parametric Model Total Variation Distance Machine Learning Efficacy + Range Query
@@
TVAE [65] Heuristically Private Variational Autoencoder Likelihood Fitness Machine Learning Efficacy
@@
CTGAN [65] Heuristically Private Generative Adversarial Network Likelihood Fitness Machine Learning Efficacy
@@
PATE-GAN [30] Differentially Private Generative Adversarial Network - Machine Learning Efficacy
@@
TabDDPM [32] Differentially Private Diffusion Model Correlations Machine Learning Efficacy
@@
TableDiffusion [60] Differentially Private Diffusion Model Kolmogorov-Smirnov Test Machine Learning Efficacy
@@
GReaT [7] Heuristically Private Large Language Model - Machine Learning Efficacy
@@
Since synthetic data is proposed to protect the privacy of real data, privacy should be an indispensable part of the evaluation. While many differentially private (DP) synthesizers [69, 39, 70, 40] have been proposed with provable privacy guarantees, they often suffer from significant utility loss [58, 20, 27]. Therefore, a parallel line of research focuses on what we call heuristically private (HP) synthesizers [48, 71, 32, 68, 7, 31], which do not integrate DP for model learning. These methods typically rely on similarity-based metrics to empirically evaluate privacy risks [^1] These metrics are also widely used in industry, see [19] for detailed analysis.. The underlying rationale is that the closer the synthetic data is to real data, the higher the risk of information leakage. However, the meaning of these metrics has not been sufficiently examined. Ironically, recent attacks [19] utilize the DCR (one of the most popular similarity-based privacy metrics, detailed in Section 4.1) to successfully recover a large proportion of outliers in real data, casting doubt on the effectiveness and reliability of current privacy measurements.
@@
Another problem of existing evaluations is the prevalent reliance on default hyperparameter settings for training synthesizers [57, 7, 68, 31]. This approach can result in suboptimal performance of synthesizers and skewed comparisons, leading to misleading conclusions (we summarize them in Appendix E). The root of this problem lies in the fact that many synthesizers do not provide details about hyperparameter selections, posing a challenge in conducting fair comparisons among diverse synthesizers on an equal footing.
@@
The above concerns motivate us to design a principled evaluation framework for data synthesis, aimed at elucidating the current advancements in this field. We believe that principled and systematic assessments are crucial: not only can they help practitioners choose from a large set of synthesizers, but they can also aid in understanding the strengths and weaknesses of each type of method, making progress toward developing better synthesis algorithms.
1.1 Our Contributions
@@
In this paper, we examine, characterize, and critique the commonly used evaluation metrics, and propose a principled framework as well as a set of new metrics for data synthesis evaluation. Our assessments unfold along three main axes:
  • Fidelity. We identify faithfulness and universality as two key properties for fidelity metrics in Section 2.1 and find that none of the existing fidelity metrics can satisfy both. To address this gap, we introduce a new fidelity metric based on the Wasserstein distance, which provides a unified way to evaluate the numerical, discrete, and mixture attributes under the same criteria.
@@
  • Privacy. We address the inadequacy of current similarity-based privacy metrics and membership inference (MI) attacks and propose the novel membership disclosure score to directly gauge the disclosure risk of synthesizers without relying on membership inference (MI) attacks. Our measurement is aligned with the principles of DP and only requires black-box access to synthesizers.
@@
  • Utility. We advocate using two tasks for assessing the utility of synthetic data: machine learning prediction and range (point) query. To eliminate the inconsistent performance caused by the choice of different machine learning models, we introduce a new and robust metric to quantify the distribution shift of synthetic data.
@@
SynMeter. We implement a systematic evaluation framework called SynMeter to support the assessment of data synthesis algorithms with the proposed metrics. SynMeter differs from previous approaches in that, instead of relying on default hyperparameters of different synthesizers, it incorporates a model tuning phase. SynMeter combines the proposed metrics to create a novel and unified objective for model tuning, which eases hyperparameter selection and consistently improves the quality of synthetic data for all methods. With a modular design, SynMeter can easily integrate any additional synthesizers and datasets by implementing new functional codes to the relevant modules. Our code is publicly available [^2] https://github.com/zealscott/SynMeter, facilitating researchers to leverage existing pipelines to tune, assess, or benchmark (new) data synthesis algorithms.
@@
Main Findings. Extensive experiments have been conducted on a wide range of state-of-the-art heuristically private (HP) and differentially private (DP) synthesis algorithms (list in Table 1) over 12 real-world datasets. Our evaluation finds that no single synthesis algorithm emerges as the definitive leader. Synthesizers struggle to balance competitive performance with robust privacy measures, exhibiting a pronounced performance disparity between HP and DP synthesis algorithms. Specifically, we have the following interesting findings:
@@
  • Diffusion models are surprisingly good at synthesizing tabular data. Although they were originally introduced for image generation, our evaluation indicates their expressive capability of synthesizing highly authentic tabular data. For instance, the diffusion-based synthesizer (i.e., TabDDPM) surpasses its counterparts in terms of both fidelity and utility. However, it suffers from significant membership privacy risks, and directly applying differential privacy to it would negate its advantages, highlighting a critical challenge for further exploration.
@@
  • Statistical methods are still competitive synthesizers. State-of-the-art statistical methods consistently outperform all deep generative synthesizers when DP is required. Impressively, even in heuristically private settings, these statistical approaches still maintain competitive performance over many sophisticated synthesizers like TVAE and GReaT.
  • Disparate impacts of privacy budget on DP synthesizers: We observe that statistical methods like PGM can achieve good performance even with a small privacy budget (e.g., $\epsilon=0.5)$. However, deep generative synthesizers generally require a much larger privacy budget (e.g., $\epsilon=8)$ to yield comparable results.
@@
  • Despite CTGAN being widely recognized as the strong baseline in data synthesis, a closer examination using proposed metrics revealed that it struggles to learn marginal distributions during training, which leads to unsatisfactory results on complex tabular data.
@@
Related Studies. Several studies [58, 27, 3, 9] have been conducted to benchmark tabular synthesis algorithms. However, they all focus on DP synthesizers, neglecting extensively studied HP synthesizers. In particular, they did not include the newly proposed deep generative synthesizers using diffusion models or large language models [32, 7, 31], which we found to be very promising in our studies. Furthermore, they directly use existing metrics for comparisons, while we identify the limitations of these metrics and propose a new and more principled assessment for tabular data synthesis.
@@
Roadmap. We first propose a principled framework for tabular data synthesis evaluation in Section 2. Then from Section 3 to Section 5, we discuss existing evaluation methods and propose our metrics from fidelity, privacy and utility aspects. In Section 6 we introduce a unified tuning objective as well as our framework implementation. The experimental results are shown in Section 7. Finally, we summarize our findings and conclude the paper in Section 8.
2 A Framework for Data Synthesis Evaluation
In this section, we introduce a framework for evaluating the effectiveness of synthesizers.
2.1 Evaluation Criteria
@@
A synthesizer $\mathsf{A}$ takes as input a dataset $\mathcal{D}$, which is assumed to be sampled i.i.d. from some (often unknown) underlying data distribution $\mathbb{D}$, and outputs a generator $\mathsf{G}$. $\mathsf{G}$ on input an integer $n$ outputs a dataset $\mathcal{S}$ with $n$ data instances. We consider three classes of desirable properties for data synthesis: Fidelity, Privacy, and Utility.
@@
Fidelity. We want the distribution for the synthetic data instances to be close to $\mathbb{D}$. Since $\mathbb{D}$ is often unknown, fidelity is measured by the distance between the input dataset $\mathcal{D}$ and the synthesized dataset $\mathcal{S}$. If one partitions the input dataset $\mathcal{D}$ into a training set $\mathcal{D}_{\text{train}}$ and a testing set $\mathcal{D}_{\text{test}}$, one can measure fidelity as closeness to either $\mathcal{D}_{\text{train}}$ or $\mathcal{D}_{\text{test}}$. Given the heterogeneity and high-dimensional nature of tabular data, assessing the distribution similarity can be challenging. We use the following desiderata for fidelity metrics:
  • Faithfulness. An effective fidelity metric should faithfully reveal the difference between the distributions of synthetic data and real data.
  • Universality. To tackle the heterogeneity of tabular data, the metric should accommodate both numerical and categorical attributions, and avoid reliance on a plethora of ad-hoc measures for different data types.
@@
Privacy. Using synthetic data is usually motivated by the desire to protect the input dataset. Without this consideration, one should share the input dataset to achieve the best fidelity and utility. For privacy, differential privacy [16] is generally considered to be the golden standard. Some synthesizers are designed to satisfy DP, either by directly perturbing data statistics or adding noises during training. However, satisfying DP under reasonable parameters may result in poor performance; heuristically private (HP) synthesizers do not satisfy DP, and aim to protect privacy empirically. We thus need empirical metrics for privacy evaluations, and the following requirements should be considered:
  • A privacy metric should reflect the foundational concept of privacy, i.e., the ability to protect every training instance against membership disclosure, as it is well understood that DP is equivalent to membership privacy when membership of different instances are independent [35]. Heuristic approaches [71] such as the closeness to real data deviate from this goal and violate this requirement.
  • A reliable privacy metric should be able to differentiate the membership disclosure risks among various synthesizers. While membership inference (MI) attacks against classifiers have been extensively studied, MI attacks against synthesizers either rely on unrealistic settings [57] or are too weak for evaluations [28]. Using MI attacks to evaluate the privacy of synthesizers is thus problematic.
We note that privacy metrics are also useful for evaluating DP synthesizers, as the privacy analysis may not be tight and the degrees of tightness vary based on the algorithms. Two algorithms with the same DP parameters $\epsilon,\delta$ may be offering different levels of privacy protection (detailed in Section 7.6).
Utility. Sometimes, when using synthesizers, one know the specific tasks the synthesized datasets will be used for, and desires that the synthesized dataset $\mathcal{S}$ is useful for tasks at hand, which may be achievable without high fidelity. We consider the following requirements for utility metrics that measure effectiveness of synthesized datasets for specific tasks.
  • The tasks chosen for evaluating synthetic data should be both representative and well-defined.
  • The performance on these tasks should be consistent and reliable. Since some tasks use auxiliary evaluators (e.g., machine learning models), it is essential to ensure that the choice of these evaluators does not affect the evaluations.
2.2 Evaluation Pipelines

### v2

@@ Abstract.
Data synthesis has been advocated as an important approach for utilizing data while protecting data privacy. A large number of tabular data synthesis algorithms (which we call synthesizers) have been proposed. Some synthesizers satisfy Differential Privacy, while others aim to provide privacy in a heuristic fashion. A comprehensive understanding of the strengths and weaknesses of these synthesizers remains elusive due to drawbacks in evaluation metrics and missing head-to-head comparisons of newly developed synthesizers that take advantage of diffusion models and large language models with state-of-the-art marginal-based synthesizers.
@@
In this paper, we present a systematic evaluation framework for assessing tabular data synthesis algorithms. Specifically, we examine and critique existing evaluation metrics, and introduce a set of new metrics in terms of fidelity, privacy, and utility to address their limitations. Based on the proposed metrics, we also devise a unified objective for tuning, which can consistently improve the quality of synthetic data for all methods. We conducted extensive evaluations of 8 different types of synthesizers on 12 real-world datasets and identified some interesting findings, which offer new directions for privacy-preserving data synthesis.
Privacy-preserving data synthesis, evaluation protocols
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Management of Data; June 22-27, 2025; Berlin, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy
1. Introduction
@@
Data-driven decision-making has emerged as the prevailing approach to advance science, industrial applications, and governance, creating the necessity to share and publish tabular data. At the same time, growing concerns about the privacy breaches caused by data disclosure call for data publishing approaches that preserve privacy. One increasingly advocated and adopted approach to reduce privacy risks while sharing data is to release synthetic data. Ideally, synthetic data can effectively fit any data processing workflow designed for the original data without privacy concerns. Data synthesis initiatives have been promoted not only by the research community (Tao et al., 2021; Jordon et al., 2021) but also among non-profit organizations (OECD, 2023; UN, 2023) and government agencies (Benedetto et al., 2018; NIST, 2018, 2020).
@@
In this paper, we study data synthesis algorithms for tabular data, which we call synthesizers. In recent years, a plethora of synthesizers have been proposed, which can be roughly categorized into two groups: statistical and deep generative. Statistical synthesizers (Zhang et al., 2017; McKenna et al., 2019, 2021; Zhang et al., 2021) use low-order marginals to create synthetic datasets that match real distributions. They were the best-performing algorithms in NIST competitions (NIST, 2018, 2020). Deep generative synthesizers (Park et al., 2018; Jordon et al., 2018; Xu et al., 2019; Kotelnikov et al., 2023; Borisov et al., 2023), on the other hand, learn the data distribution from real data and generate synthetic instances by sampling from the distribution. With the recent development in deep generative models (e.g., diffusion models (Ho et al., 2020) and large language models (Vaswani et al., 2017; Radford et al., 2019)), new synthesizers are proposed (Kotelnikov et al., 2023; Borisov et al., 2023; Kim et al., 2022; Zhang et al., 2023b) to extend these successes to the realm of tabular data synthesis.
@@
While recent state-of-the-art approaches achieve compelling results in synthesizing authentic data, a comprehensive understanding of the strengths and weaknesses of different synthesizers remains elusive. We are unaware of any head-to-head comparison of the best statistical synthesizers (McKenna et al., 2019, 2021; Zhang et al., 2021) against the best deep generative synthesizers (Kotelnikov et al., 2023; Borisov et al., 2023; Kim et al., 2022).
@@
In addition, there is a lack of principled and widely accepted evaluation metrics. It is known that evaluating synthesizers is inherently difficult (Theis et al., 2016; Chundawat et al., 2022; Alaa et al., 2022; Jordon et al., 2021), and qualitative evaluation of tabular data through visual inspection is also infeasible. Table 1 shows that existing papers use different evaluation metrics, resulting in substantially disparate conclusions.
@@
Table 1. Main evaluation protocols for synthesizers in the literature. “-” means that the synthesizer does not consider this property. “ ✓ ” denotes that the model is viewed as the state-of-the-art method by many related studies (McKenna et al., 2022; Ganev et al., 2022; Ouyang et al., 2023; Tao et al., 2021). “ ✓   ” denotes that the model is frequently used in recent papers(Kotelnikov et al., 2023; Lee et al., 2023; Kim et al., 2022) for comparison as a state-of-the-art method.
@@
PGM (McKenna et al., 2021, 2019) Differentially Private Probabilistic Graphical Model Total Variation Distance Range Query
@@
PrivSyn (Zhang et al., 2021) Differentially Private Non-parametric Model Total Variation Distance Machine Learning Efficacy + Range Query
@@
TVAE (Xu et al., 2019) Heuristically Private Variational Autoencoder Likelihood Fitness Machine Learning Efficacy
@@
CTGAN (Xu et al., 2019) Heuristically Private Generative Adversarial Network Likelihood Fitness Machine Learning Efficacy
@@
PATE-GAN (Jordon et al., 2018) Differentially Private Generative Adversarial Network - Machine Learning Efficacy
@@
TabDDPM (Kotelnikov et al., 2023) Differentially Private Diffusion Model Correlations Machine Learning Efficacy
@@
TableDiffusion (Truda, 2023) Differentially Private Diffusion Model Kolmogorov-Smirnov Test Machine Learning Efficacy
@@
GReaT (Borisov et al., 2023) Heuristically Private Large Language Model - Machine Learning Efficacy
@@
Since synthetic data is proposed to protect the privacy of real data, privacy should be an indispensable part of the evaluation. While many differentially private (DP) synthesizers (Zhang et al., 2017; McKenna et al., 2021; Zhang et al., 2021; McKenna et al., 2022) have been proposed with provable privacy guarantees, they often suffer from significant utility loss (Tao et al., 2021; Ganev et al., 2022; Hu et al., 2024). Therefore, a parallel line of research focuses on what we call heuristically private (HP) synthesizers (Park et al., 2018; Zhao et al., 2021; Kotelnikov et al., 2023; Zhang et al., 2023b; Borisov et al., 2023; Kim et al., 2022), which do not integrate DP for model learning. These methods typically rely on similarity-based metrics to empirically evaluate privacy risks [^1] These metrics are widely used both in academia and industry as default privacy evaluation protocols, see (Ganev and De Cristofaro, 2023) for detailed analysis.. The underlying rationale is that the closer the synthetic data is to real data, the higher the risk of information leakage. However, the meaning of these metrics has not been sufficiently examined. Ironically, recent attacks (Ganev and De Cristofaro, 2023) utilize these similarity-based metrics to successfully recover a large proportion of outliers in real data, casting doubt on the effectiveness and reliability of current privacy measurements.
@@
Another problem of existing evaluations is the prevalent reliance on default hyperparameter settings for training synthesizers (Stadler et al., 2022; Borisov et al., 2023; Zhang et al., 2023b; Kim et al., 2022). This approach can result in suboptimal performance of synthesizers and skewed comparisons, leading to misleading conclusions [^2] We discuss these misleading conclusions in our technical report version.. The root of this problem lies in the fact that many synthesizers do not provide details about hyperparameter selections, posing a challenge in conducting fair comparisons among various synthesizers on an equal footing.
@@
The above concerns motivate us to design a systematic evaluation framework for data synthesis to elucidate the current advancements in this field. We believe that principled and systematic assessments are crucial: not only can they help practitioners choose from a large set of synthesizers, but they can also aid in understanding the strengths and weaknesses of each type of method, making progress toward developing better synthesis algorithms.
1.1. Our Contributions
@@
In this paper, we examine, characterize, and critique the commonly used evaluation metrics, and propose a systematic framework as well as a set of new metrics for data synthesis evaluation. Our assessments unfold along three main axes:
  • Fidelity. To address the heterogeneity and high dimensionality of tabular data, we introduce a new fidelity metric based on Wasserstein distance, providing a unified way to evaluate the numerical, discrete, and mixture marginals under the same criteria.
@@
  • Privacy. We identify the inadequacy of existing similarity-based metrics and membership inference attacks and propose a novel privacy metric called membership disclosure score to directly gauge the disclosure risk of synthesis algorithms. Our measurement aligns with the principles of differential privacy and only requires black-box access to synthesizers.
@@
  • Utility. We advocate two tasks for assessing the utility of synthesizers: machine learning prediction and range (point) query. To eliminate the inconsistent performance caused by the choice of different machine learning models, we introduce a new and robust metric to quantify the distribution shift of synthetic data.
@@
$\mathsf{SynMeter}$. We implement a systematic evaluation framework called $\mathsf{SynMeter}$ to support the assessment of data synthesis algorithms with the proposed metrics. Differing from the existing evaluations, $\mathsf{SynMeter}$ incorporates the model tuning phase, which eases hyperparameter selection and consistently improves the performance of synthesizers for fair comparison. With a modular design, $\mathsf{SynMeter}$ can easily integrate additional synthesizers and datasets by implementing new functional codes to the relevant modules. Our code is publicly available [^3] https://anonymous.4open.science/r/SynMeter, facilitating researchers to leverage existing pipelines to tune, assess, or benchmark (new) synthesis algorithms.
@@
Main Findings. Extensive experiments have been conducted on a wide range of state-of-the-art heuristically private (HP) and differentially private (DP) synthesizers (list in Table 1) over 12 real-world datasets. Our evaluation finds that no single synthesis algorithm emerges as the definitive leader. Synthesizers struggle to balance competitive performance with robust privacy measures, exhibiting a pronounced performance disparity between HP and DP synthesis algorithms. Specifically, we have the following interesting findings:
@@
  • Diffusion models are surprisingly good at synthesizing tabular data. Although they were originally introduced for image generation, our evaluation indicates their impressive capability of synthesizing highly authentic tabular data. For instance, the HP diffusion-based synthesizer (i.e., TabDDPM) surpasses its counterparts and nearly reaches the empirical upper bound in terms of both fidelity and utility. However, it suffers from significant membership privacy risks, and directly applying differential privacy to it would negate its advantages, highlighting a critical challenge for further exploration.
@@
  • Statistical methods are still competitive synthesizers. State-of-the-art statistical methods consistently outperform all deep generative synthesizers when DP is required, especially when the privacy budget is small (e.g., $\epsilon=0.5)$. Impressively, even in heuristically private settings, these statistical approaches still maintain competitive performance and empirical privacy protection over many sophisticated synthesizers like TVAE and GReaT.
@@
  • Despite CTGAN being widely recognized as the strong baseline in data synthesis, a closer examination using proposed metrics reveals that it struggles to learn marginal distributions, which leads to unsatisfactory results on complex tabular data.
@@
Related Studies. Several studies (Tao et al., 2021; Hu et al., 2024; Arnold and Neunhoeffer, 2020; Bowen and Snoke, 2019) have been conducted to benchmark tabular synthesis algorithms. However, they all focus on DP synthesizers, neglecting extensively studied HP synthesizers. In particular, they did not include the newly proposed deep generative synthesizers using diffusion models or LLM (Kotelnikov et al., 2023; Borisov et al., 2023; Kim et al., 2022), which we found to be very promising in our studies. Furthermore, they directly use existing metrics for comparisons, while we identify the limitations of these metrics and propose a new set of metrics as well as the systematic assessment for tabular data synthesis.
@@
Roadmap. We first propose a systematic framework for tabular data synthesis evaluation in Section 2. Then from Section 3 to Section 5, we discuss existing evaluation methods and propose our metrics from fidelity, privacy, and utility. In Section 6 we introduce a unified tuning objective and our framework implementation. The experimental results are shown in Section 7. We summarize our findings and offer some suggestions for practitioners in Section 8.
2. A Framework for Data Synthesis Evaluation
2.1. Evaluation Criteria
@@
A synthesizer $\mathsf{A}$ takes as input a dataset $\mathcal{D}$, which is assumed to be sampled i.i.d. from some (often unknown) underlying data distribution $\mathbb{D}$, and outputs a generator $\mathsf{G}$. Given an integer $n$, $\mathsf{G}$ outputs a dataset $\mathcal{S}$ with $n$ data instances. We consider three classes of desirable properties for synthesizers: Fidelity, Privacy, and Utility.
@@
Fidelity. As the substitute for real data, the distribution for the synthetic data instances should be close to $\mathbb{D}$. Since $\mathbb{D}$ is often unknown, fidelity is measured by the distance between the input dataset $\mathcal{D}$ and the synthetic dataset $\mathcal{S}$. If one partitions the input dataset $\mathcal{D}$ into a training set $\mathcal{D}_{\text{train}}$ and a testing set $\mathcal{D}_{\text{test}}$, one can measure fidelity as closeness to either $\mathcal{D}_{\text{train}}$ or $\mathcal{D}_{\text{test}}$.
@@
Privacy. Using synthetic data is usually motivated by the desire to protect the input dataset. Without this consideration, one may share the input dataset to achieve the best fidelity and utility. For privacy, differential privacy (Dwork, 2006) is generally considered to be the golden standard. Some synthesizers are designed to satisfy DP, either by directly perturbing data statistics or adding noises during training. However, satisfying DP under reasonable parameters may result in poor performance. Heuristically private (HP) synthesizers do not satisfy DP, and aim to protect privacy empirically. Therefore, privacy metrics are necessary to evaluate HP synthesis algorithms. In addition, privacy metrics are also useful for evaluating DP synthesizers, as the privacy analysis may not be tight and the degrees of tightness vary based on the algorithms. Two algorithms with the same DP parameters $\epsilon,\delta$ may offer different levels of privacy protection (analyzed in Section 7.5).
Utility. Synthetic data often replaces real datasets to perform specific downstream tasks. In these cases, the synthetic data may not necessarily need to preserve high fidelity, as long as it achieves good utility for these tasks. Hence, utility evaluation is useful to measure the effectiveness of synthesizers for common tasks.
2.2. Evaluation Pipelines